### v2

# Probes of Anomalous Events at LHC with Self-Organizing Maps

Abstract
We propose an Unsupervised Learning Algorithm, Self-Organizing Maps (SOM), built on a neural network architecture, for the probe of a rare top decay, mediated by Flavor Changing Neutral Current (FCNC), to charm and the Higgs boson, with the Higgs boson further decaying to a pair of b-quarks or a pair of gauge bosons ($W^{\pm}/Z$) in a boosted regime. Ideally, the particles originating from the decay of the boosted top lead to the reconstruction of a large-R jet, comprising three-prong substructures, with b- and c-tagged subjets. We have illustrated that SOM can be effectively implemented as an anomaly-finder for model-agnostic probes of such decay at the LHC. Relatively simple in formalism, this algorithm achieves commendable performance over other existing formalisms, tailored to probe this rare decay. We discuss the robustness of this algorithm, which can be successfully implemented for other BSM probes, both for model-agnostic and model-dependent searches.
1 Introduction
The Standard Model of Particle Physics (SM) has been an accomplished framework encapsulating strong and weak interactions of nature [1, 2, 3, 4]. Albeit its success, theoretical inconsistencies, and observational discrepancies motivated theories Beyond the Standard Model (BSM) [5, 6, 7]. With several BSM theories proposed [8], and with recent upgrades at the Large Hadron Collider (LHC) underway [9, 10], the identification of events governed by a specific BSM formalism over abundant SM backgrounds is becoming increasingly challenging, establishing the pursuit of advanced and effective identification strategies is a necessity.
Recent advancements in Machine Learning (ML) algorithms have progressively replaced traditional cut-based search techniques to identify boosted objects at the LHC [11, 12, 13, 14, 15]. According to the ML dialect, BSM searches with supervised learning techniques reduce to a classification problem for a model-dependent probe, where events are categorized individually as signals or backgrounds depending on their respective traits. On the contrary, model-agnostic probes with unsupervised learning techniques identify anomalies over patterns anticipated from collisions governed by the SM framework on a collective dataset. In this paper, we introduce an ML algorithm, the Self-Organizing Maps (SOM) [16] and demonstrate its effectiveness for the latter search strategy.
The landscape of ML-driven top-tagging has evolved significantly. Traditional methods [17, 18] have been surpassed by sophisticated approaches, encompassing both model-dependent and model-agnostic probes leveraging respectively, supervised and unsupervised learning strategies. For a glimpse into the diverse array of such topics, we refer the reader to [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39], and [40, 41] for a comprehensive review of the domain. However, there has been limited exploration of SOM [42] in Particle Physics Phenomenology. The scope of this work is a portrayal of SOM as an anomaly-probing prescription for model-independent searches, demonstrated over FCNC-mediated rare decay of top quark at LHC.
Self-Organizing Map resembles a planar grid, with nodes placed over the vertices. Each node represents a specific event topology. During training, SOM organizes events with similar traits together using an iterative approach, forming groups that capture events with underlying similarities together. In this paper, we explore the potential of this algorithm as an effective tool for anomaly detection, as well as an efficient identifier of a specific BSM signature. We observe that SOM possesses promising implications, associated not only with the identification of the rare decay processes over an overwhelming SM background but also with the segregation of different decay modes of heavy states with similar event topologies.
As an illustration, we probed a rare top decay, driven by a flavor changing neutral current (FCNC), to a c-quark and Higgs boson $\left(t\to cH\right)$ with suppressed branching fraction $\lesssim 0.1\%$ [43, 44] in a boosted regime ($p^{top}_{T}>$ 350 GeV), over the background dominated by the most abundant decay mode of the top $\left(t\to bW\right)$, and spurious QCD events $\left(pp\to jj\right)$ mimicking the signal topology. The decay products of the boosted top are highly collimated, enveloped by a large-R jet comprising a three-prong substructure at LHC. Kinematic features specific to the signal $\left(t\to cH\right)$, irreducible background $\left(t\to bW\right)$, and the QCD di-jet background, are mapped on the planar grid which, through iterative adjustment of weights attributed on the nodes, leads to the identification of individual decay modes as clusters appearing on the grids. This probe of rare top decay mediated by FCNC using SOM achieves a signal significance of $\sim 2\sigma$ for $\mathcal{BR}(t\to cH)$ = 0.05% at 3 ab-1 over SM backgrounds, a fair leap as compared to our previous endeavor [45], where we presented a simple ML model for the same probe that outperformed existing top-tagging methodologies [40], tailored to probe the said decay. Finally, we provide an interpretation for each cluster boundary formed on the SOM grid, drawing insights into the relevant features underlying the event topologies of the three different processes under consideration.
The approach followed for this work is model-independent and robust; similar prescriptions can be formulated for BSM probes involving boosted composite topologies. With anticipated operations of future particle colliders in higher energy frontiers, we find the analysis of boosted objects pertinent, with potentials for witnessing New Physics breakthroughs in the near future.
This paper has been organized as follows: Section 2 provides an overview of the SOM algorithm, Section 3 outlines the event generation, object reconstruction, and input feature selection processes, Section 4 discusses the results obtained from the training, providing insights into the patterns reflected on the trained weight vectors on the grid, Section 5 articulates the optimal signal region for the FCNC top decay process ($t\to cH$), and Section 6 provides the concluding remarks, future directives, and highlights of the entire exploration.
2 Self Organizing Maps: The Algorithm
The Self Organizing Maps (SOM) is an Unsupervised Machine Learning algorithm that generates a lower-dimensional representation of a higher-dimensional dataset, simultaneously preserving the topological structure of the data. Through Competitive Learning, SOM learns to pin down the position of every training data-point (simulated events in this context) on the topographic map of neurons. Consequently, dimensionality reduction makes SOM a perfect tool for data visualization.
An SOM, unlike generic Deep Learning architectures, does not require a target output for training. Instead, where the weights assigned to a given node match the input vector (a specific event), the weights on the nodes present in the neighborhood of the Best Matching Unit (BMU) are selectively optimized to resemble the data point, subsequently capturing the underlying traits of the class (signal or different backgrounds) the input vector belongs to.
From an initial distribution of random weights and over subsequent iterations, SOM eventually converges into a map of stable zones, through an iterative process of determining the node best resembling a given input event, labeled the Best Matching Unit (BMU), thereby updating the weights of neurons locally encircling the BMU, in accordance with the neighborhood function and respective differences from the input data.
The updated weight vector for a specific node is given by:
$W_{v}(t+1)=W_{v}(t)+\eta(t)\cdot\phi(u,v,t)\cdot[(X(t)-W_{v}(t))],$ (1)
where, $v$ are the coordinates of the said node, $u$ are the coordinates of the BMU, $W_{v}(t)$ is the weight vector of the said node at time step $t$, $X(t)$ is the input feature presented at time step $t$, $\eta(t)$ is time-dependent learning rate, and $\phi(u,v,t)$ is the Neighborhood Distance Function for the BMU (at $u$) and the concerned node (at $v$) at the time step $t$. The learning rate $\eta(t)$ follows an exponential damping given as,
$\eta(t)=\eta(0)\cdot\exp{(-t/\lambda)},$ (2)
with $\eta(0)$ being the initial learning rate and $\lambda$ being the learning decay rate. The Neighborhood Distance Function follows a Gaussian distribution pattern. It decays rapidly with the distance between the said node and the BMU, with the decay rate impeded by a time-dependent factor mimicking the variance of the Gaussian distribution. Hence, Eqn. 1 implies that the nodes closest to the BMU are dominantly aligned with the input feature, with the effect of the alignment attenuating on moving away from the BMU for a given time step. Neighborhood Distance Function is given by,
$\phi(u,v,t)=\exp{\left(\frac{-[d(u,v)]^{2}}{2[\sigma(t)]^{2}}\right)},$ (3)
with $d(u,v)$ being the distance metric between the coordinates $u$ and $v$, and $\sigma(t)$ is the attenuating factor at time step t following exponential distribution that reads,
$\sigma(t)=\sigma(0)\exp{(-t/\beta)},$ (4)
where $\sigma(0)$ is the initial attenuating factor and $\beta$ is the attenuation decay rate.
We employ Batch Self-Organizing Map (BSOM) to ensure efficient training over the entire dataset at once, potentially leading to faster model convergence and a more stable cluster boundary. The input features are scaled using the MinMaxScaler normalization scheme so that every value for a given feature ranges between 0 and 1, and all input features receive equal importance in the model training. Unlike traditional SOMs that update weights after considering each data point, BSOMs update weights after processing a batch of data. Treating the entire training dataset as a single batch impedes the model convergence, and varying the batch-size we find that batches of 500 events yield faster convergences and stable outputs. Prioritizing the simplicity of the model, we choose linearly decreasing values for the learning rate ($\eta$) and the attenuating factor ($\sigma$) with each batch. Starting from 1, we varied $\eta$ to 0, and $\sigma$ from 20 to 1. We use the implementation of the SOM algorithm provided in the GitHub repository [46], with the necessary modifications mentioned above.
3 Event Generation and Reconstructed Objects
This section provides an overview of the action plan followed for the preparation of input data required for training the self-organizing map architecture. We sketch below the steps undertaken for event generations, object reconstructions, and feature selections:
  • Generation of Monte Carlo Samples:
    We have considered two dominant SM processes: $(i)$ associated production of a top-quark with a $W$-boson, with $W$-boson undergoing leptonic decay and the top decaying to a bottom quark and a hadronically decaying $W$ boson, and $(ii)$ production of di-jet QCD events for training the network. In order to achieve the boosted regime, we simulate the processes putting a generation level cut of 350 GeV on the minimum transverse momentum of the top quark and QCD parton (quark/gluon). We use MadGraph (version 3.4.1) [47] to generate hard scattering events at the leading-order (LO) in $\alpha_{s}$ followed by the desired decays using MadSpin. The collision energy has been considered to be 13 TeV, and both the renormalization and factorization scales are fixed at $H_{T}$/2, where $H_{T}$ denotes the cumulative transverse energy of the colliding partons. We consider the Parton Distribution Function (PDF) given by NNPDF23-lo-as-0130-qed [48]. The generated events are passed through PYTHIA8 (version 8.310) [49] to account for the initial and final state radiations, parton showering and hadronization. The HepMC [50] output obtained from PYTHIA8 is further fed into Delphes (version 3.5.0) [51] to consider the detector effects, using the CMS card.
    For the signal events, we follow a model-independent, generic framework [52] leading to the rare top decay $\left(t\to cH\right)$, and use FeynRules [53, 54] to generate the UFO output, which is provided to MadGraph to generate hard scattering events.
  • Object Reconstruction:
    We use FastJet [55, 56] with anti-kT algorithm [57] for the reconstruction of jets, with jet radius set at $R=1.0$, and optimized based on the truth-level event information to capture the decay products of the top quark, aligning the jet-axis along the direction of the top. To mitigate contamination from underlying events (UE) and multiple parton interactions (MPI), jet trimming [58] is applied. This involves reclustering the enveloped products into subjets with jet radius 0.3 using kT algorithm, and removing the subjets with momentum fraction less than 3% of the large-R jet momentum. Reconstructed large-R jets with $p_{T}>$ 400 GeV and mass $\in$ [140, 200] GeV are finally selected for further analysis; these are the candidate top-jets.
    For the identification of flavor-tagged subjets, we identify b- and c-hadrons within the large-R jet from particle-level information. A subjet is labeled b- or c-tagged if the angular separation $(\Delta R)$ between the corresponding hadron and the subjet-axis is less than the subjet radius. Detector effects are considered for the flavor-tagging by imposing b- and c-tagging efficiencies to be 70% and 40% respectively, and using a $p_{T}$-dependent formula for mistagging probabilities [59, 60]. To prevent multiple tagging, b-tagging is performed first on $p_{T}$-ordered subjets, and the tagged subjets are excluded from subsequent flavor-tagging. Subjets failed to be flavor-tagged are labeled as light jets.
  • Selection of Jet Substructure based Observables as Input Features:
    We have followed the approach outlined in [45] for the consideration of relevant jet substructure based observables, presenting a comprehensive list of the features used for SOM training in Table 1.
    Input features for SOM
    $p_{T}^{b-jet}$ $\Delta E_{T}^{b-jet1}$ $\Delta E_{T}^{c-jet}$ $\tau_{21}^{1}$ $\Delta X_{cj}$ $\Delta X_{bj}$ $N_{c-jet}$
    $p_{T}^{c-jet}$ $\Delta E_{T}^{b-jet2}$ $\tau_{32}^{1}$ $\Delta X_{cb}$ $N_{b-jet}$ $N_{Ljet}$
    Table 1: The set of input variables used for training SOM.
    We briefly sketch an outline of few important subjet based features below, referring the interested readers to [45] for further details:
      N-subjettiness ($\tau^{\beta}_{N}$): N-subjettiness characterizes the substructure of boosted jets [61]. Considering N hypothetical subjet axes, N-subjettiness is determined by summing up the angular distances of jet constituents with respect to their respective closest subjet axes. Essentially, it indicates how well a jet aligns with N subjet-structure hypothesis. To establish the prominence of the existence of three subjet axes within the large-R jets of the signal (or the irreducible background) over two, and two subjet axes over one (for QCD di-jet background), we analyze the ratios of N-subjettiness measurements, specifically measuring $\tau_{N}^{\beta}/\tau_{N-1}^{\beta}$, labeled $\tau_{N(N-1)}^{\beta}$, for respectively $N=3,~{}2$, considering the angular weightage exponent $\beta=1$. These ratios enables the isolation of the QCD multijets from top jet samples.
      Subjet Multiplicities & Transverse Momenta: Multiplicities of light and flavor-tagged subjets $\left(N_{b-jet},~{}N_{c-jet}\right)$ along with the transverse momentum of leading flavored subjet $\left(p_{T}^{b-jet},~{}p_{T}^{c-jet}\right)$ are crucial for the identification of signal and different SM backgrounds.
      Energy- and Mass-Fraction Information: The energy repatriation of the large-R jet is evaluated by the fraction of its total energy stored in its constituent subjets. For a given subjet, this is measured as, $\Delta E_{T}^{subjet}=\frac{E_{subjet}}{E_{topjet}}$. We further estimate the composition of large-R jet, with an observable measuring the fractional mass retained after removing the invariant mass of specific subjet combinations (indexed by $i,~{}j$): $\Delta X_{ij}=\left(1-\frac{M_{ij}}{M_{topjet}}\right)$. Accurate identification of light jets is crucial for this analysis, and is achieved through a $\chi^{2}$-analysis detailed in [45].
4 Results
Having trained an SOM with QCD di-jet and SM top decay $(t\to bW)$ samples, we discuss our observations in this section.
4.1 Training on Data and the U-Matrix:
Figure 1: Left: The trained SOM, displaying clusters formed from top decay and QCD multijet events. Black lines indicate regions with no data points, separating distinct clusters. Right: The U-Matrix visualization of the trained SOM, revealing a vertical boundary dividing the map in a roughly 1:2 ratio, with the larger right-hand section showing further sub-clusters.
A Self-Organizing Map (SOM) with a $60\times 70$ grid is trained over 500,000 events, each event represented by thirteen features listed in Table 1. The vector weights over each node are initialized at random. Taking into account each event, the comparison between the event feature vector and the vector weights is drawn on each node, the Best Matching Unit is identified, and the weight vectors are modified according to Equation 1. Through multiple iterations on the training dataset, clusters appear and distinct cluster boundaries emerge, as shown in the left panel of Figure 1. The cluster boundaries, which are the darker regions of the grid, represent areas without data points. They effectively envelope the brighter regions, which capture events with similar underlying traits. Moving forward, each cluster would unfold a unique predominant event topology encountered in the training dataset. The intensity of the color reflects the density of the events on the Map, scaled by the total number of events used to train the grid.
We determine the strength of the identified cluster boundaries with a U-Matrix. The elements of U-Matrix are calculated for every node by measuring the average Euclidean distance between the weight vector on the concerned node and that on its neighboring nodes. This visualization allows us to identify the most significant cluster separations. The U-Matrix that emerges from our trained SOM is presented in the right panel of Figure 1. The U-Matrix separates the grid with a vertical boundary leading to 2 major clusters, along with 5 smaller, more tightly defined sub-clusters, within the cluster at right. The detailed analysis of these boundaries and the event populations within each cluster is presented in the following sections.
4.2 Robustness of the cluster boundaries
We have taken a step further to assess the topographic error for the obtained map, which determines the fraction of events for which the Best Matching Unit and second Best Matching Unit are adjacent to each other. Figure 2 illustrates the evolution of the topographic error in the first 50 iterations. The error starts at 3.5% and stabilizes around 3.3%, indicating an effective model convergence.
Figure 2: Topographic error after training the map with QCD multijet and SM top decay sample for 50 iterations.
To further assess the quality of the clustering, we employ two evaluation metrics: the Silhouette score [62] and the Davies-Bouldin Index (DBI) [63].
  • Silhouette score: The Silhouette score measures how well an event fits within its assigned cluster compared to other clusters. For a given event e, is defined as:
    $S=\frac{b-a}{\max(b,a)}$
    where, ‘$a$’ denotes the mean distance between e and all other events within its own cluster, and ‘$b$’ is the mean distance between e and all events in its nearest-neighboring cluster. The average value of $S$ for all events gives the average Silhouette score. This score ranges between $[-1,1]$, where positive values indicate moderate to good clustering and negative values suggest poor clustering. A score close to zero implies that the events are globally populated in the vicinity of the cluster boundaries.
  • Davies-Bouldin Index (DBI): This metric is used to assess the quality of the clustering. A lower DBI value signifies the presence of compact, well-separated clusters. DBI is given by,
    $\text{DBI}=\frac{1}{N}\sum_{i=1}^{N}\max_{i\neq j}\left(\frac{S_{i}+S_{j}}{D_{ij}}\right)$
    where, $N$ is the total number of clusters, $S_{i}$ is the mean intra-cluster distance for the $i^{th}$ cluster (a measure of compactness), and $D_{ij}$ denotes the distance between the centroids of the clusters $i$ and $j$ (a measure of separation). The condition $\max_{i\neq j}$ allows the identification of the least compact cluster $j$ and close to the cluster $i$ (worst case similarity).
We implement these two metrics from sklearn library [64] to evaluate the clustering performance achieved after training. We obtain a Silhouette score of 0.46, which indicates that the trained model achieves moderate clustering quality, with reasonably well-defined clusters. A DBI score of 1.04 supports this claim, suggesting fair cluster compactness and separation. These metrics confirm satisfactory model performance on the training dataset. For cross-verification, we compared these metrics evaluated for the trained SOM with those obtained from the k-means clustering. The result is in close agreement and demonstrates consistency and robustness.
4.3 Analysis of the Weight Planes and Clusters
Figure 3: The prominent clusters as seen from Figure 1. The clusters are assigned with specific numbers for convenience. The events inside each of these clusters are then identified for further analysis.
As stated in Section 2, SOM provides a unique data visualization tool to analyze the underlying structure of the training data, without any prior knowledge of its topography. Basically, SOM clusters the data points (events) by identifying the underlying structure and grouping similar data points together. Therefore, the cluster boundary, functionally, resembles a gradient where the neighboring nodes of the grid represent progressively different data traits. Based on the representation of the U-matrix in the right panel of Figure 1, we identify seven (7) clusters leveled as C-1 to C-7, as shown in Figure 3, to proceed with further analysis.
An insight into the weight planes would provide a better understanding of the clusters and their boundaries. The “weight”, in this context, carries an entirely different meaning than its conventional sense for Artificial Neural Networks (ANN). For ANNs, the values corresponding to each input feature are multiplied by a weight, and their total contribution is measured by passing them as an argument for an activation function. For SOMs, the weights are assigned to the output nodes, without any activation function. Through adjustments of these weights, SOM creates insights into the feature space. Hence, to gain a deeper understanding of how these features shape different collision outcomes, we now discuss their corresponding weight planes.
Figure 4: Weight vector plane corresponding to the features, $N_{b-jet}$, $N_{c-jet}$, $N_{Ljet}$, $\Delta E_{T}^{b-jet1}$, $\Delta E_{T}^{b-jet2}$, $\Delta E_{T}^{c-jet}$, respectively. The unified effect of these weights is what we see in the U-Matrix in Figure 1.
In the upper panel of Figure 4, we display the weight planes corresponding to different jet multiplicities, namely, the number of b and c-tagged jets and light jets. The clearly visible boundaries in the weight planes for $N_{b-jet}$ and $N_{c-jet}$ highlight the role these two features play in the clustering process. For example, the dark region on the left side of the weight plane of $N_{b-jet}$ represents the grids that correspond to events having negligible number of b-tagged jets - a well-known feature of the QCD di-jet events. The map of the weights of $N_{c-jet}$ supports this claim, as these events are also expected to have a very small number of c-tagged jets. Events with non-zero values of both b- and c-tagged jet multiplicities are mapped to the lower-right corner of the map. As expected, light jet multiplicity did not play much of a role as both the QCD and top events contain a large number of light jets. Moving to the lower panel of Figure 4, where we show the map of weight planes for three other decisive features, namely $\Delta E_{T}^{b-jet1}$, $\Delta E_{T}^{b-jet2}$, and $\Delta E_{T}^{c-jet}$. As seen in the figures, these features help to understand the energy profile among the flavor-tagged subjets. These observables are also mapped with visible decision boundaries in the weight planes, significantly contributing to the different clusters discussed in Figure 3.
Figure 5: Weight plane corresponding to the features, $\Delta X_{bj}$ , $\Delta X_{cj}$, and $\Delta X_{cb}$ respectively.
The presence of a W boson within the large R top jet motivated us to include some observables in terms of the invariant mass of the flavor-tagged subjets, namely $\Delta X_{bj}$, $\Delta X_{cj}$, and $\Delta X_{cb}$. As evident in Figure 5, along with the multiplicities of the jet and their energy profile, these features greatly help clustering events. In fact, the brightest boundaries in the U-Matrix are primarily determined using the multiplicities of light-jet, b-jet, and c-jet, along with the energy and mass distributions among these subjets.
The significant observations obtained through the weight plane analysis are as follows:
  • The leftmost cluster (viz., C-1) includes mostly the events having high light-jet multiplicity with (almost) no b- or c-tagged jets. Therefore, this cluster certainly corresponds to the QCD multijet events.
  • The right subclusters (other than C-1) have events with one or more b-tagged jets and light-jets. Additionally, the events mapped at the bottom right corner (viz., C-2 to C-5), in addition, possess non-zero c-tagged jets. Therefore, the right-hand cluster corresponds to the events from the top decay. Clusters with 2 or more b-tagged jets would include events originating from non-standard decay of the top quark.
  • There exist at least two clusters, namely C-4 and C-6, which contain events that have two or more b-tagged jets along with at least one c-tagged jet. These events are not expected to originate from the SM decay of the top quarks and, therefore, could be considered as clusters having “anomalous” events.
4.4 Clustering efficiency with Test dataset
To assess the performance of the SOM training, we test the model with three kinds of samples,
  1. 1.
    Test dataset1: $pp\to tW$ events with the top following the dominant decay mode producing b- and c-quarks in the final state ($t\to bW^{+}$ with $W^{+}\to c\bar{s}$ and the corresponding charge-conjugated process). The $W$-boson produced in association with the top quark decays leptonically.
  2. 2.
    Test dataset2A/2B/2C: $pp\to tW$ events with top quark decaying through an FCNC decay, $t\to cH$ followed by $H\to b\bar{b}$/$H\to WW^{*}$/$H\to ZZ^{*}$ decay. The $W$-boson produced in association with the top quark decays leptonically.
  3. 3.
    Test dataset3A/3B: The most dominant channel, $pp\to t\bar{t}$, with one of the tops decaying leptonically111 The simulation methodology for the $t\bar{t}$ events, including the software suites, is identical to that described in Section 3. $(t\to bW,W\to\ell\nu_{\ell})$, and the other top undergoes $t\to bW$ decay or $t\to cH$ decay with $W$/$H$ decaying hadronically.
A summary of the test data sets is shown in Table 2.
Dataset Process considered
Test dataset1 $pp\to Wt~{}(p_{T}^{t}>350~{}GeV),\;W\to\mu~{}\nu_{\mu};$
(SM) $t\to bW,\;W\to c\;s$
Test dataset2A $pp\to Wt~{}(p_{T}^{t}>350~{}GeV),\;W\to\mu~{}\nu_{\mu};$
(BSM) $t\to cH,\;H\to b\;\bar{b}$
Test dataset2B $pp\to Wt~{}(p_{T}^{t}>500~{}GeV),\;W\to\mu~{}\nu_{\mu};$
(BSM) $t\to cH,\;H\to WW^{*},\;W\to jj$
Test dataset2C $pp\to Wt~{}(p_{T}^{t}>500~{}GeV),\;W\to\mu~{}\nu_{\mu};$
(BSM) $t\to cH,\;H\to ZZ^{*},\;Z\to jj$
[…]
(SM) $\bar{t}\to bW,\;W\to j\;j$
[…]
(BSM) $\bar{t}\to cH,\;H\to b\;\bar{b}$
Table 2: The details of the three types of dataset considered in testing the trained SOM.
For validation, we calculate the efficiencies for independent (unseen) datasets of the SM top and QCD di-jet events. Almost 90% of the QCD di-jet samples accumulated in the left clusters (e.g., C-1 and C-2) of the map, while 90% of the SM top samples are assigned to the right clusters (C-3 to C-7) as portrayed in the first two rows of Table 3. We next proceed with the analysis using the test datasets. Figure 6 shows the clustering of the two test datasets, namely Test dataset1 and Test dataset2A, in respectively left and right figures. The weight plane analysis suggests that Test dataset1 is dominantly clustered in the right half of the SOM grid, as these regions correspond to non-zero multiplicities of both b- and c-tagged jets. On the other hand, for Test dataset2A, which has higher multiplicities of b-tagged jet, should be clustered primarily in C-4 and C-6. Both figures confirm our expectation, and therefore motivate us to calculate the cluster efficiencies, defined as the number of events captured in a given cluster over the total number of events analyzed, for each cluster for a given test dataset. We select 100,000 events for each test dataset that satisfy the pre-selection cuts and trigger requirements, and tabulate the efficiencies in Table 3. The higher efficiencies in C-5, C-6 and C-7 for Test dataset1, and C-4 and C-6 for Test dataset2A, clearly reflect the performance of the training algorithm. The phase space of the other two test datasets corresponding to BSM physics, namely Test dataset2B and Test dataset2C, are expected to have strong overlap with the QCD multijet events. As the numbers in the last two rows of Table 3 reveal, these events are indeed clustered in C-1 and C-2, as expected. Furthermore, we observe that $t\bar{t}$ events, like single top quark events, follow expected patterns. Test dataset3A characterized by the dominant top decay mode, are principally captured in cluster C-7. In contrast, events for Test dataset3B are distributed among the clusters C-4, C-6, and C-7, revealing a more diverse event topology, as shown in the last two rows of Table 3. Based on these observations, it is reasonable to expect that a combined analysis of $tW$ and $t\bar{t}$ production processes would improve the achievable limit on the top FCNC decay.
Figure 6: Left: The projection of the events corresponding to the two test datasets - "Test dataset1" (left panel) and "Test dataset2A" (right panel). We refer the main text for more details.
Process C-1 C-2 C-3 C-4 C-5 C-6 C-7
SM top 5.45% 4.83% 2.51% 0.21% 12.19% 6.81% 67.15%
QCD di-jet 82.84% 7.69% 0.23% 0.03% 0.47% 1.91% 6.26%
Test dataset1 4.48% 5.39% 5.04% 0.42% 22.48% 11.02% 49.62%
Test dataset2A 2.19% 0.88% 3.65% 12.45% 7.67% 61.06% 7.99%
Test dataset2B 46.93% 36.78% 0.36% 0.035% 2.67% 0.51% 11.08%
Test dataset2C 49.06% 35.37% 0.45% 0.049% 2.22% 0.67% 10.96%
Test dataset3A 4.1% 4.86% 2.44% 0.1% 11.5% 5.5% 70.77%
Test dataset3B 0.9% 1% 3.5% 12.22% 7.9% 56.42% 15.27%
Table 3: Cluster efficiencies, defined as the number of events clustered in the given cluster over the total number of events analyzed, corresponding to each cluster for a given test dataset. The dominant contribution(s) of each dataset towards the clusters are expressed in bold letters.
5 BSM Reach: Event Selection to Signal Significances
[…]
For those events which pass the pre-selection cuts, we identify three signal regions (SR) based on the flavor tagging information among the subjets of the large-R jet. The signal regions are primarily motivated by the understanding of the different clusters obtained after training the SOM and their characteristics in terms of the high-level objects. Additionally, for a comparative analysis, the SRs are also defined keeping in mind the resolved (non-boosted) analysis performed by the ATLAS Collaboration to find the top FCNC decay at the 13 TeV run of LHC.
Figure 7: The invariant mass distribution of the two b-tagged jets for SR-I (left) and SR-II (right) without the invariant mass cut.
The first signal region (SR-I) is almost the same as the ATLAS study, with the only difference being that the requirements are made at the subjet level. We extend the selection strategy of SR-I to SR-II, accommodating the further requirement of c-tagged jets within the large-R jet. It is needless to say that SR-II is expected to be the most efficient selection strategy in probing the top FCNC decay we are interested in. Finally, we also study the sensitivity of a relatively relaxed criterion to the multiplicity of b-tagged jets within the large R jet by introducing the third signal region, SR-III. Note that, for both SR-I and SR-II, the presence of at least two b-tagged jets enables us to calculate the invariant mass of the two b-tagged jets, and compare the reconstructed mass with the Higgs boson mass. In Figure 7, we show the distribution of the invariant mass mentioned above for the SM Top samples, Test dataset2A, Test dataset3A and Test dataset3B that satisfies the criteria for SR-I and SR-II without the invariant mass cut. For SR-I and SR-II, therefore, we further demand that the invariant mass of the two leading b-tagged jets satisfy the following: $M_{bb}=[110,140]~{}{\rm GeV}$. The details of the signal regions are listed in Table 4.
SR-I nSubjet $\geq$ 3,   $N_{b-jet}\geq$ 2,   $M_{bb}=[110,140]~{}{\rm GeV}$
SR-II nSubjet $\geq$ 3,   $N_{b-jet}\geq$ 2,   $N_{c-jet}>$ 0,   $M_{bb}=[110,140]~{}{\rm GeV}$
SR-III nSubjet $\geq$ 3,   $N_{b-jet}>$ 0,  $N_{c-jet}>$ 0
Table 4: Signal regions defined to probe the top FCNC decay. Here the observables $nSubjet$, $N_{b-jet}$ and $N_{c-jet}$ denote the total number of subjets, number of b-tagged jets and number of c-tagged jets within the large-R jet, respectively. The invariant mass of the two leading b-tagged jets is denoted by the quantity $M_{bb}$.
As noted previously, the FCNC interactions between the top quark and the Higgs boson via the $t-c-H$ coupling are essential in our quest for BSM physics. To evaluate the feasibility of detecting this exotic mode in the boosted regime using the SOM algorithm, we compute the signal significance, defined as $\mathcal{S}=\sqrt{2\left[(S+B)\ln\left(1+\frac{S}{B}\right)-S\right]}$ [65] where S and B represent the number of signal and background events that meet the jet selection criteria and satisfy the trigger/pre-selection requirement. Before we proceed to calculate the signal significances for the three signal regions, it is important to estimate the cluster efficiencies for each of the SRs. In Table 5, we show the leading order (in $\alpha_{s}$) production cross section for all the processes considered here with $\sqrt{s}=13~{}{\rm TeV}$. The last column in Table 5 estimates the expected number of events at the 13 TeV run of LHC with integrated luminosity $\mathcal{L}=3000~{}fb^{-1}$. It is important to note that even though QCD di-jet events seem to overshadow signal events as seen in $3^{rd}$ column of $2^{nd}$ row, the number of events that eventually satisfy the trigger and jet selection cuts, and contribute to the signal regions defined in Table 4 is negligible.
Process Cross-section ($fb$) $\mathcal{BR}$ $\times\mathcal{BR}\times\mathcal{L}$
SM top 435 $0.11\times 1.0\times 0.68$ 9.76×$\times 10^{4}$
QCD di-jet $3.91\times 10^{6}$ 1 1.17×$\times 10^{10}$
Test dataset2A 435 $0.11\times 1.0\times 0.6$ 8.61×$\times 10^{4}$
Test dataset2B 72.3 $0.11\times 1.0\times 0.215\times 0.674^{2}$ 2330
Test dataset2C 72.3 $0.11\times 1.0\times 0.0264\times 0.547^{2}$ 188.5
Test dataset3A 7879 $0.11\times 1.0\times 0.68\times 1.0$ $1.77\times 10^{6}$
Test dataset3B 7879 $0.11\times 1.0\times 0.6\times 1.0$ $1.56\times 10^{6}$
Table 5: The production cross section and branching ratio for the different processes considered in this analysis. The last column denotes the expected number of events at the 13 TeV run of LHC with integrated luminosity $\mathcal{L}=3000~{}fb^{-1}$.
In order to calculate the signal significance, we estimate the number of events captured in each cluster region for the BSM test datasets that pass the trigger and jet selection cuts. In Tables 6 - 8, we show the effective number of events that are clustered in each of those clusters that satisfy all the selection cuts for SR-I to SR-III. The notation S2A, S2B, S2C and S3B denotes the effective number of events for the four test datasets, viz., Test datasets 2A, 2B, 2C and 3B, respectively.
C-1 C-2 C-3 C-4 C-5 C-6 C-7
B1 - - - 55.6 8.8 1673.8 -
B2 - - - 796.5 194.7 25346.4 -
B - - - 852.1 203.5 27020.2 -
S2A - - 0.9 6426.5 539.8 16229.9 -
S2B - - - 0.33 - - -
S2C - - - 0.025 - - -
S3B - - 78 111196.8 9141.6 282375.6 -
S - - 78.9 117623.7 9681.4 298605.5 -
Table 6: The effective number of events clustered in each cluster that satisfy all the selection cuts for SR-I. The notation ‘-’ indicates that no events exist corresponding to the given cluster.
C-1 C-2 C-3 C-4 C-5 C-6 C-7
B1 - - - 55.6 8.8 - -
B2 - - - 796.5 194.7 - -
B - - - 852.1 203.5 - -
S2A - - 0.9 6426.5 539.8 - -
S2B - - - 0.33 - - -
S2C - - - 0.025 - - -
S3B - - 78 111196.8 9141.6 - -
S - - 78.9 117623.7 9681.4 - -
Table 7: The effective number of events clustered in each cluster that satisfy all the selection cuts for SR-II.
C-1 C-2 C-3 C-4 C-5 C-6 C-7
B1 - - 237.2 218.6 11795 - -
B2 - - 3062.1 2637.3 203992.5 - -
B - - 3299.3 2855.9 215787.5 - -
S2A - - 1195.1 10840 6750.2 - -
S2B - - 1.7 0.9 62 - -
S2C - - 0.3 0.1 4.2 - -
S3B - - 19968 181755.6 124066.8 - -
S - - 21165.1 192596.6 130883.2 - -
Table 8: The effective number of events clustered in each cluster that satisfy all the selection cuts for SR-III.
Considering an example, we define S2A as,
$S2A=\sigma\times BR\times\mathcal{L}\times\epsilon_{cl},$ (5)
[…]
Signal Region S B $\mathcal{BR}(t\to cH)$ $\mathcal{S}\,$
0.1% 2.5
SR-I 425989.5 28075.8 0.05% 1.3
0.03% 0.8
0.1% 3.8
SR-II 127384 1055.6 0.05% 1.9
0.03% 1.2
0.1% 0.7
SR-III 344644.9 221942.7 0.05% 0.4
0.03% 0.2
Table 9: Signal significances for three SRs defined based on different choices of flavor tagging at the subjet level.
[…]
6 Summary and Outlook
Probes of processes involving rare decays driven by Flavor-Changing Neutral Current (FCNC) offer a powerful avenue for exploring Physics beyond the Standard Model. The rare decay, $t\to cH$, considered in this article, offers sensitive probe of potential New Physics interactions, in particular, the $t-c-H$ coupling. In contrast to the supervised learning methodology presented in [45], we present a complementary, model-independent analysis based on unsupervised learning technique, targeting the boosted regime of this decay. With judicious choices of jet-substructure based observables along with efficient flavor tagging of subjets, similar to the approach followed in [45], we achieve an enhanced signal significance over our previous analysis [45], which outperformed the performances of advanced ML-driven approaches, viz., GNN-based LorentzNet and Transformer-based classifiers.
In [45], we argued that the judicious choice of jet-substructure based observables provided an edge over the GNN-based models which train on the 4-vectors of the large-R jet constituents. With the jet substructures of the signal $\left(t\to cH,~{}H\to b\bar{b}\right)$, bearing close resemblance to those of the irreducible background $\left(t\to bW,~{}W\to c\bar{s}\right)$, this schema achieved relatively lower significance. The transformer-based approaches implement attention mechanism, which determines the relevance of input features for proper identification signals from the background, thereby allowing a dense deep-learning classifier selectively focus on relevant features suppressing the irrelevant ones, consequently enhancing the accuracy and efficiency of the signal recognition. We observed that a simple boosting algorithm with decision tree stumps trained over judiciously chosen jet-substructure-based features, nevertheless, exhibited higher efficiency. This indicates that our choice of input features made the feature space so fragmented that the dominance of a specific class (signal or different SM backgrounds) in a given subspace becomes prominent.
The Self Organizing Maps, on the contrary, follows a different approach. The nodes, placed at the intersection of its grids, attempt to mimic the feature vector for each input data-point (event topologies) iteratively. The weights are adjusted so that the best matching unit (BMU) and the nodes in its neighborhood have their respective weights aligned with the matched feature vector. With training over each iteration (and updating weights assigned to each node), and subsequently over every batch (and subsequent updation of model parameters), the model captures the latent features specific to signal and different backgrounds. Finally, the grids end up forming clusters, with every node in each cluster possessing a weight vector resembling the underlying features that distinguish different event classes. As discussed in Section 5, each cluster predominantly captures events belonging to a specific Test dataset (see, Tables 6 to 8). The clusters formed on SOM grid are the planar projection of the feature space of the events, in a way, that allows events with similar underlying topology captured in a single cluster.
Training SOM with SM top quark and QCD di-jet samples with assessment of topographic error in clustering, we observe that the emerged map fairly preserves the data topography within the grid. Further testing of the trained model with FCNC-driven top decay to a charm quark and a Higgs boson ($t\to cH$) demonstrates remarkable performance in identification of the anomalous regions for the training samples. We define three signal regions (SRs) based on the three-prong substructure of the candidate top jet, and their kinematics. For SR-II, our model achieves a significance of $\sim 2\sigma$ for $\mathcal{BR}(t\to cH)$ = 0.05%, establishing the model performance for segregating rare FCNC-mediated decays from overwhelming SM backgrounds. We emphasize that a possible explanation behind an enhanced performance of SOM over the existing approaches is embedded in the working principle of the algorithm, which learns to mimic the event features, leading to better capturing of patterns specific to the signal and different backgrounds in distinct clusters.
We have considered the implementation of SOM as an unsupervised anomaly detector in Section 4, achieving fairly substantial significance despite vanishingly small $t\to cH$ branching fraction over an overwhelming SM background. We commenced with training the SOM architecture with SM backgrounds and used the trained model to probe the rare top-decay as anomalous events over the identified SM patterns. SOM can also be tailored for supervised learning, upon training the grids with the signal events and the SM backgrounds, thereby estimating the clusters and the event density in each cluster, using SOM as a model-dependent probe of the rare top-decay.
SOMs, since its inception, has outgrown as tool for clustering and dimensionality reduction. As evoked above, it has been adapted for supervised learning, enabling enhanced model performance, simultaneously preserving the topography of the training dataset. This versatility makes SOM a potential tool for BSM probes. Further improvements can be explored with alternative distance metrics, viz., the Mahalanobis distance, or the cosine similarity, in order to obtain more distinct clusters of signals and backgrounds on Kohonen plane. Adaptive frameworks, such as, Growing Self-Organizing Map (GSOM), which dynamically adapts its structure during training, by adding nodes as required based on data complexity, could further optimize the signal-to-background ratio. SOMs and its extensions are therefore, potential tools for future phenomenological explorations.
7 Acknowledgments
AC acknowledges the financial support from the Department of Science and Technology, Government of India, under Grant No. IFA18-PH 224 (INSPIRE Faculty Award). The authors express their gratitude to the organizers of the ML4HEP-2023 conference held at ICTS Bangalore, India, for the initial development of this project. AC thanks Biplob Bhattacherjee for useful comments, and SC acknowledges Benjamin Fuks for insightful discussions at the Iwate Collider School-2025 at Iwate, Japan. The authors also thank Tousik Samui and Dharitree Bezboruah for their input in the initial stages of the work.